# Conditional entropy and information of quantum processes

### v1

@@ IV.1 Recipe for deriving information measures of quantum channels
The completely mixing map $\mathcal{R}_{A\rightarrow B}$ is a linear completely positive map which replaces input trace-class operators with a scalar times the identity operator; in particular, its action on the trace-class operators $X\in\mathrm{L}(A)$,
$\mathcal{R}_{A\rightarrow B}(X):=\mathrm{tr}(X)\mathbbm{1}_{B}.$ (29)
@@
The generalized entropy $\mathbf{S}[\mathcal{N}]$ of a point-to-point channel $\mathcal{N}_{A^{\prime}\to A}$ is defined as the generalized channel divergence between the quantum channel $\mathcal{N}$ and the completely mixing map $\mathcal{R}$ [19, 23],
@@
$S(A|B)_{\rho}$ $=S(AB)_{\rho}-S(A)_{\rho}=-D(\rho_{AB}\|\mathbbm{1}_{A}\otimes\rho_{B})=\inf_{\omega\in{\mathrm{S}t}(B)}D(\rho_{AB}\|\mathbbm{1}_{A}\otimes\omega_{B}),$ (33)
@@
The von Neumann entropy, conditional entropy, and mutual information of states $\rho_{AB}$ (Eqs. (32)-(34)) have been extended (generalized) in the literature using appropriate generalized state divergences $\mathbf{D}$ in the following way (cf. [7]).
@@
    $-D[\mathcal{N}_{A^{\prime}B^{\prime}\to AB}\|\mathcal{R}_{A\to A}\circ\mathcal{N}_{A^{\prime}B^{\prime}\to AB}].$ (38)
  • Candidate B. Taking the form $\mathcal{R}_{A^{\prime}\to A}\otimes\mathcal{M}_{B^{\prime}\to B}$, where $\mathcal{M}\in{\mathrm{Q}}(B^{\prime},B)$, to be a possible channelization of $\mathbbm{1}_{A}\otimes\omega_{B}$ as $\omega\in{\mathrm{S}t}(B)$. $\mathbbm{1}_{A}$ is substituted by $\mathcal{R}_{A^{\prime}\to A}$ and state $\omega_{B}$ is substituted by channel $\mathcal{M}_{B^{\prime}\to B}$. That would give the following potential candidate for the conditional entropy of $\mathcal{N}_{A"B^{\prime}\to AB}$,
    $-\inf_{\mathcal{M}\in{\mathrm{Q}}(B^{\prime},B)}D[\mathcal{N}_{A^{\prime}B^{\prime}\to AB}\|\mathcal{R}_{A\to A}\otimes\mathcal{M}_{B^{\prime}\to B}].$ (39)
@@
We first begin to discuss with Candidate B as our preferred definition for the conditional entropy of the channel. We do so because the form of Candidate B is generalizable to the definition for mutual information of channel. We will, in the latter sections, see that Candidate B seems operationally preferable to Candidate A even if we don’t have a generalization of the form to multipartite mutual information in mind.
@@ IV.2 Definitions and properties
where the infimum is over all channels $\mathcal{M}_{B^{\prime}\to B}$. In particular, if we optimize (supremum) $\mathbf{D}_{\mathrm{A}ccess}$ over the set $\mathrm{E}_{e}=\{\rho_{RA^{\prime}B^{\prime}}:\operatorname{tr}[\widehat{H}_{RAB}\rho_{RA^{\prime}B^{\prime}}]\leq e\in\mathbb{R}^{+}\}$ of input states $\rho_{RA^{\prime}B^{\prime}}$ with bounded energy ($\leq e\in(0,\infty)$), then $\mathbf{S}_{\mathrm{E}_{e}}[A|B]_{\mathcal{N}}$, with $\mathrm{Acess}=\mathrm{E}_{e}$, is called energy-constraint generalized conditional entropy.
@@
where $P[\mathcal{K},\mathcal{L}]$ is purified channel distance function [31] between $\mathcal{K},\mathcal{L}\in{\mathrm{Q}}(A^{\prime},A)$,
@@ V.1 Conditional von Neumann entropy
$S[A|B]_{\mathcal{N}}:=\inf_{\mathcal{M}\in{\mathrm{Q}}(B^{\prime},B)}D[\mathcal{N}_{A^{\prime}B^{\prime}\to AB}\|\mathcal{R}_{A^{\prime}\to A}\otimes\mathcal{M}_{B^{\prime}\to B}],$ (58)
@@
We know from [47, Lemma 1] that, for an arbitrary state $\rho_{AB}$ and $\sigma\in{\mathrm{P}os}(A)$,
$\inf_{\omega\in{\mathrm{S}t}(B)}D(\rho_{AB}\|\sigma_{A}\otimes\omega_{B})=D(\rho_{AB}\|\sigma_{A}\otimes\rho_{B}).$ (60)
@@
$-S[A|B]_{\mathcal{N}}$ $=\inf_{\mathcal{M}\in{\mathrm{Q}}(B^{\prime},Q)}D[\mathcal{N}_{A^{\prime}B^{\prime}\to AB}\|\mathcal{R}_{A^{\prime}\to A}\otimes\mathcal{M}_{B^{\prime}\to B}]$
$=\inf_{\mathcal{M}\in{\mathrm{Q}}(B^{\prime},Q)}\sup_{\outerproduct{\phi}{\phi}\in{\mathrm{S}t}(RA^{\prime}B^{\prime})}D(\mathcal{N}(\phi_{RA^{\prime}B^{\prime}})\|\mathcal{R}\otimes\mathcal{M}(\phi_{RA^{\prime}B^{\prime}}))$
$=\inf_{\mathcal{M}\in{\mathrm{Q}}(B^{\prime},Q)}\sup_{\outerproduct{\phi}{\phi}\in{\mathrm{S}t}(RA^{\prime}B^{\prime})}D(\mathcal{N}(\phi_{RA^{\prime}B^{\prime}})\|\mathbbm{1}_{A}\otimes\mathcal{M}(\phi_{RB^{\prime}}))$
@@
Consider a reduced channel $\mathcal{T}^{\mathcal{N}}_{B^{\prime}\to B}$ obtained from a bipartite tele-covariant channel $\mathcal{N}_{A^{\prime}B^{\prime}\to AB}$ [39, Definition 4], when $|A^{\prime}|,|A|<+\infty$, and the maximally mixed state $\pi_{A^{\prime}}$,
$\mathcal{T}^{\mathcal{N}}_{B^{\prime}\to B}(\cdot):=\operatorname{tr}_{A}\circ\mathcal{N}_{A^{\prime}B^{\prime}\to AB}(\pi_{A^{\prime}}\otimes\cdot),$ (62)
then the bipartite channels $\frac{1}{|A|}\mathcal{R}_{A^{\prime}\to A}\otimes\mathcal{T}^{\mathcal{N}}_{B^{\prime}\to B}$ and $\mathcal{N}_{A^{\prime}B^{\prime}\to AB}$ are always jointly tele-covariant (see [39, Remark 1] and [48, Definition 5]).
Consider a reduced channel $\mathcal{T}^{\mathcal{N}}_{B^{\prime}C^{\prime}\to B}$ obtained from a tele-covariant tripartite channel $\mathcal{N}_{A^{\prime}B^{\prime}C^{\prime}\to ABC}$ [9, Definition 24],
@@
From Lemma 2 and [9, Remark 5], it directly follows that the bipartite channel $\frac{1}{|A|}\mathcal{R}_{A^{\prime}\to A}\otimes\mathcal{T}^{\mathcal{N}}_{B^{\prime}\to B}$ is jointly tele-covariant with the channel $\mathcal{N}_{A^{\prime}B^{\prime}\to AB}$, where $\mathcal{T}^{\mathcal{N}}$ is defined in Lemma 2. Then, $-S[A|B]_{\mathcal{N}}\leq{D}[\mathcal{N}\|\mathcal{R}\otimes\mathcal{T}^{\mathcal{N}}]={D}(\Phi^{\mathcal{N}}\|\frac{1}{|A^{\prime}|}\mathbbm{1}_{R_{A}A}\otimes\Phi^{\mathcal{T}^{\mathcal{N}}})=\log|A^{\prime}|-S(R_{A}A|R_{B}B)_{\Phi^{\mathcal{N}}}$, where we have used the result [31, Corollary II.5] for the first equality. We now prove the converse part to conclude the proof. For a quantum channel $\mathcal{M}_{B^{\prime}\to B}$ that gives infimum for ${D}[\mathcal{N}_{A^{\prime}B^{\prime}\to AB}|\mathcal{R}_{A^{\prime}\to A}\otimes\mathcal{M}_{B^{\prime}\to B}]$, we have
@@
Inequality (68) is a direct consequence of [19, Proposition 6]. From [19, Corollary 8], we can say that $S[AB]_{\mathcal{N}}=-\log|A^{\prime}||B^{\prime}|$ if and only if $\mathcal{N}_{A^{\prime}B^{\prime}\to AB}$ is an isometric channel. ∎
@@
We now calculate the von Neumann conditional entropy for some tele-covariant two-qubit quantum channels of practical interest in quantum computing and information processing tasks.
Figure 1: In this figure, we plot the von Neumann conditional entropies for partial SWAP, partial CNOT, partially noisy SWAP, and partially noisy CNOT gates with respect to the corresponding channels’ parameter $p$. The horizontal line at $S[A|B]_{\mathcal{S}^{P}}=-1$ is for the two-qubit identity gate $\operatorname{id}_{A^{\prime}\to A}\otimes\operatorname{id}_{B^{\prime}\to B}$. We can notice from the plot that the partial SWAP, partial CNOT, and noisy SWAP gates have $S[A|B]_{\mathcal{N}}<-1$ for all $p\in[0,1)$.
Partial SWAP and CNOT operation.— We consider the two-qubit noiseless partial SWAP and CNOT operations [50, 51, 52] given by $\mathcal{S}^{p}_{A^{\prime}B^{\prime}\to AB}(\cdot)=U^{p}_{\mathrm{SWAP}}(\cdot)U^{p}_{\mathrm{SWAP}}$, where ${U}^{p}_{\mathrm{SWAP}}=\sqrt{p}\mathbbm{1}+\iota\sqrt{1-p}\mathrm{SWAP}$ and $\mathcal{C}^{p}_{A^{\prime}B^{\prime}\to AB}(\cdot)=U^{p}_{\mathrm{CNOT}}(\cdot)U^{p}_{\mathrm{CNOT}}$ where ${U}^{p}_{\mathrm{CNOT}}=\sqrt{p}\mathbbm{1}+\iota\sqrt{1-p}\mathrm{CNOT}$, respectively for $p\in[0,1]$.
Following Corollary 2, the von Neumann conditional entropy of the two-qubit noiseless partial SWAP $\mathcal{S}^{p}_{A^{\prime}B^{\prime}\rightarrow AB}$ and CNOT operation $\mathcal{C}^{p}_{A^{\prime}B^{\prime}\rightarrow AB}$
$S[A|B]_{\mathcal{S}^{P}}$ $=-S(R_{B}B)_{\Phi^{\mathcal{S}^{p}}}-\log|A^{\prime}|=3\left(\frac{1-p}{4}\log{\frac{1-p}{4}}\right)+\left(\frac{1+3p}{4}\log{\frac{1+3p}{4}}\right)-1,$ (73)
$S[A|B]_{\mathcal{C}^{P}}$ $=-S(R_{B}B)_{\Phi^{\mathcal{C}^{p}}}-\log|A^{\prime}|=\frac{1-\sqrt{p}}{2}\log{\frac{1-\sqrt{p}}{2}}+\frac{1+\sqrt{p}}{2}\log{\frac{1+\sqrt{p}}{2}}-1.$ (74)
For $p=0$, $\mathcal{S}^{0}$ is the two-qubit SWAP operation and $S[A|B]_{\mathcal{S}^{0}}=-3$ (cf. Remark 2). For $p=1$, $\mathcal{S}^{1}$ is the two-qubit identity operation and $S[A|B]_{\mathcal{S}^{1}}=-1$ (cf. Theorem A4).
Partially noisy SWAP and CNOT gates.— We consider the two-qubit SWAP and CNOT gates mixed with the two-qubit identity gate. The two-qubit partially noisy SWAP and CNOT gates are given by $\mathcal{N}^{p,\mathrm{SWAP}}_{A^{\prime}B^{\prime}\to AB}(\cdot)=p\operatorname{id}(\cdot)+(1-p)\mathrm{SWAP}(\cdot)\mathrm{SWAP}^{\dagger}$ and $\mathcal{N}^{p,\mathrm{CNOT}}_{A^{\prime}B^{\prime}\to AB}(\cdot)=p\operatorname{id}(\cdot)+(1-p)\mathrm{CNOT}(\cdot)\mathrm{CNOT}^{\dagger}$, respectively, where $p\in[0,1]$. Following Corollary 2, the von Neumann conditional entropies of $\mathcal{N}^{p,\mathrm{SWAP}}_{A^{\prime}B^{\prime}\to AB}$ and $\mathcal{N}^{p,\mathrm{CNOT}}_{A^{\prime}B^{\prime}\to AB}$ are
@@
[…]
where $a_{\pm}=\frac{1}{2}\left(1\pm\sqrt{3p^{2}-3p+1}\right)$.

### v2

@@ IV.1 Recipe for deriving information measures of quantum channels
The completely mixing map $\mathcal{R}_{A^{\prime}\rightarrow A}$ is a linear completely positive map which replaces input trace-class operators with a scalar times the identity operator; in particular, its action on the trace-class operators $X\in\mathrm{L}(A^{\prime})$ is
$\mathcal{R}_{A^{\prime}\rightarrow A}(X):=\mathrm{tr}(X)\mathbbm{1}_{A}.$ (31)
The Choi operator of the completely mixing map $\mathcal{R}_{A^{\prime}\rightarrow A}$ is $\Gamma_{RA}^{\mathcal{R}}=\mathbbm{1}_{R}\otimes\mathbbm{1}_{A}$, where $R\simeq A^{\prime}$.
@@
The generalized entropy $\mathbf{S}[\mathcal{N}]$ of a point-to-point channel $\mathcal{N}_{A^{\prime}\to A}$ is defined as the generalized channel divergence between the quantum channel $\mathcal{N}\in{\mathrm{Q}}(A^{\prime},A)$ and the completely mixing map $\mathcal{R}\in{\mathrm{C}P}(A^{\prime},A)$ [20, 24],
@@
$S(A|B)_{\rho}$ $=S(AB)_{\rho}-S(B)_{\rho}=-D(\rho_{AB}\|\mathbbm{1}_{A}\otimes\rho_{B})=-\inf_{\omega\in{\mathrm{S}t}(B)}D(\rho_{AB}\|\mathbbm{1}_{A}\otimes\omega_{B}),$ (35)
@@
We note that the negative conditional entropy $-S(A|B)_{\rho}$ of a bipartite state $\rho_{AB}$ is equal to the coherent information $I(A\rangle B)_{\rho}$, a information metric that is of significance in the context of quantum communication [48, 5].
The von Neumann entropy, conditional entropy, and mutual information of states $\rho_{AB}$ (Eqs. (34)-(36)) have been extended (generalized) in the literature using appropriate generalized state divergences $\mathbf{D}$ in the following way (e.g., [8, 27, 46]).
@@
    $S^{\not\to}[A|B]_{\mathcal{N}}:=-D[\mathcal{N}_{A^{\prime}B^{\prime}\to AB}\|\mathcal{R}_{A\to A}\circ\mathcal{N}_{A^{\prime}B^{\prime}\to AB}].$ (40)
  • Candidate B. Taking the form $\mathcal{R}_{A^{\prime}\to A}\otimes\mathcal{M}_{B^{\prime}\to B}$, where $\mathcal{M}\in{\mathrm{Q}}(B^{\prime},B)$, to be a possible channelization of $\mathbbm{1}_{A}\otimes\omega_{B}$ as $\omega\in{\mathrm{S}t}(B)$. $\mathbbm{1}_{A}$ is substituted by $\mathcal{R}_{A^{\prime}\to A}$ and state $\omega_{B}$ is substituted by channel $\mathcal{M}_{B^{\prime}\to B}$. This would give the following potential candidate for the conditional entropy of $\mathcal{N}_{A"B^{\prime}\to AB}$,
    $S[A|B]_{\mathcal{N}}=-\inf_{\mathcal{M}\in{\mathrm{Q}}(B^{\prime},B)}D[\mathcal{N}_{A^{\prime}B^{\prime}\to AB}\|\mathcal{R}_{A^{\prime}\to A}\otimes\mathcal{M}_{B^{\prime}\to B}].$ (41)
@@
We first begin to discuss with Candidate B as our preferred definition for the conditional entropy of the channel. We do so because the form of Candidate B is generalizable to the definition for mutual information of channel. We will, in the latter sections, see that Candidate B seems operationally preferable to Candidate A in general.
@@ IV.2 Definitions and properties
where the infimum is over all channels $\mathcal{M}_{B^{\prime}\to B}$. In particular, if we optimize (supremum) $\mathbf{D}_{\mathrm{A}ccess}$ over the set $\mathrm{E}_{e}=\{\rho_{RA^{\prime}B^{\prime}}:\operatorname{tr}[\widehat{H}_{RAB}\rho_{RA^{\prime}B^{\prime}}]\leq e\in\mathbb{R}^{+}\}$ of input states $\rho_{RA^{\prime}B^{\prime}}$ with bounded energy ($\leq e\in(0,\infty)$), then $\mathbf{S}_{\mathrm{E}_{e}}[A|B]_{\mathcal{N}}$, with $\mathrm{Acess}=\mathrm{E}_{e}$, is called the energy-constraint generalized conditional entropy.
@@
where $P[\mathcal{K},\mathcal{L}]$ is purified channel distance [51] between $\mathcal{K},\mathcal{L}\in{\mathrm{Q}}(A^{\prime},A)$,
@@ V.1 Conditional von Neumann entropy
$S[A|B]_{\mathcal{N}}:=-\inf_{\mathcal{M}\in{\mathrm{Q}}(B^{\prime},B)}D[\mathcal{N}_{A^{\prime}B^{\prime}\to AB}\|\mathcal{R}_{A^{\prime}\to A}\otimes\mathcal{M}_{B^{\prime}\to B}],$ (60)
@@
$=\inf_{\mathcal{M}\in{\mathrm{Q}}(B^{\prime},B)}\sup_{\outerproduct{\phi}{\phi}\in{\mathrm{S}t}(RA^{\prime}B^{\prime})}D(\mathcal{N}(\phi_{RA^{\prime}B^{\prime}})\|\mathcal{R}\otimes\mathcal{M}(\phi_{RA^{\prime}B^{\prime}}))$
$=\inf_{\mathcal{M}\in{\mathrm{Q}}(B^{\prime},B)}\sup_{\outerproduct{\phi}{\phi}\in{\mathrm{S}t}(RA^{\prime}B^{\prime})}D(\mathcal{N}(\phi_{RA^{\prime}B^{\prime}})\|\mathbbm{1}_{A}\otimes\mathcal{M}(\phi_{RB^{\prime}}))$
@@
Here the inequality holds due to [52, Lemma 1] which states that, for an arbitrary state $\rho_{AB}$ and $\sigma\in{\mathrm{P}os}(A)$, we have $\inf_{\omega\in{\mathrm{S}t}(B)}D(\rho_{AB}\|\sigma_{A}\otimes\omega_{B})=D(\rho_{AB}\|\sigma_{A}\otimes\rho_{B})$. ∎
@@
Consider a reduced channel $\mathcal{T}^{\mathcal{N}}_{B^{\prime}\to B}$ obtained from a bipartite tele-covariant channel $\mathcal{N}_{A^{\prime}B^{\prime}\to AB}$ [42, Definition 4], when $|A^{\prime}|,|A|<+\infty$, and the maximally mixed state $\pi_{A^{\prime}}$,
$\mathcal{T}^{\mathcal{N}}_{B^{\prime}\to B}(\cdot):=\operatorname{tr}_{A}\circ\mathcal{N}_{A^{\prime}B^{\prime}\to AB}\big{(}\pi_{A^{\prime}}\otimes(\cdot)\big{)},$ (63)
then the bipartite channels $\frac{1}{|A|}\mathcal{R}_{A^{\prime}\to A}\otimes\mathcal{T}^{\mathcal{N}}_{B^{\prime}\to B}$ and $\mathcal{N}_{A^{\prime}B^{\prime}\to AB}$ are always jointly tele-covariant (see [42, Remark 1] and [53, Definition 5]).
Consider a reduced channel $\mathcal{T}^{\mathcal{N}}_{B^{\prime}C^{\prime}\to B}$ obtained from a tele-covariant tripartite channel $\mathcal{N}_{A^{\prime}B^{\prime}C^{\prime}\to ABC}$ [10, Definition 24],
@@
From Lemma 2 and [10, Remark 5], it directly follows that the bipartite channel $\frac{1}{|A|}\mathcal{R}_{A^{\prime}\to A}\otimes\mathcal{T}^{\mathcal{N}}_{B^{\prime}\to B}$ is jointly tele-covariant with the channel $\mathcal{N}_{A^{\prime}B^{\prime}\to AB}$, where $\mathcal{T}^{\mathcal{N}}$ is defined in Lemma 2. Then, $-S[A|B]_{\mathcal{N}}\leq{D}[\mathcal{N}\|\mathcal{R}\otimes\mathcal{T}^{\mathcal{N}}]={D}(\Phi^{\mathcal{N}}\|\frac{1}{|A^{\prime}|}\mathbbm{1}_{R_{A}A}\otimes\Phi^{\mathcal{T}^{\mathcal{N}}})=\log|A^{\prime}|-S(R_{A}A|R_{B}B)_{\Phi^{\mathcal{N}}}$, where we have used the result [34, Corollary II.5] for the first equality. We now prove the converse part to conclude the proof. For a quantum channel $\mathcal{M}_{B^{\prime}\to B}$ that gives infimum for ${D}[\mathcal{N}_{A^{\prime}B^{\prime}\to AB}|\mathcal{R}_{A^{\prime}\to A}\otimes\mathcal{M}_{B^{\prime}\to B}]$, we have
@@
Inequality (69) is a direct consequence of [20, Proposition 6]. From [20, Corollary 8], we can say that $S[AB]_{\mathcal{N}}=-\log|A^{\prime}||B^{\prime}|$ if and only if $\mathcal{N}_{A^{\prime}B^{\prime}\to AB}$ is an isometric channel. ∎
@@
We now analytically and numerically calculate the von Neumann conditional entropy using Theorem 4 for some tele-covariant quantum channels of practical interest in quantum computing and information processing tasks.
Figure 1: In this figure, we plot the von Neumann conditional entropies $S[A|B]_{\mathcal{N}^{p,U}}$, where $U\in\{{\mathrm{S}WAP,CNOT}\}$, for noisy SWAP and CNOT operations with respect to the channels’ parameter $p$, given by Eqs. (75) and (76). The horizontal line at $S[A|B]_{\mathcal{N}}=-1$ is for the two-qubit noiseless identity gate.
@@
$S[A|B]_{\mathcal{N}^{p,{\mathrm{S}WAP}(d)}}=-\left(p+\frac{1-p}{d^{4}}\right)\log\left(p+\frac{1-p}{d^{4}}\right)-(d^{4}-1)\left(\frac{1-p}{d^{4}}\log\frac{1-p}{d^{4}}\right)-3\log d,$ (74)
where $|A^{\prime}|=|B^{\prime}|=|A|=|B|=d$ and ~$\widetilde{\mathcal{R}}_{A^{\prime}B^{\prime}\to AB}$ is the completely mixing channel, i.e., ~$\widetilde{\mathcal{R}}_{A^{\prime}B^{\prime}\to AB}=\widetilde{\mathcal{R}}_{A^{\prime}\to A}\otimes\widetilde{\mathcal{R}}_{B^{\prime}\to B}$.
SWAP and CNOT gates with white noise.— We now consider the two-qubit SWAP and CNOT gates mixed with the completely mixing channel ~$\widetilde{\mathcal{R}}_{A^{\prime}B^{\prime}\to AB}$ and $|A^{\prime}|=|B^{\prime}|=|A|=|B|=2$. The two-qubit noisy SWAP and CNOT gates are given by $\mathcal{N}^{p,\mathrm{SWAP}(2)}_{A^{\prime}B^{\prime}\to AB}(\cdot)$ and $\mathcal{N}^{p,\mathrm{CNOT}}_{A^{\prime}B^{\prime}\to AB}(\cdot)=(1-p)\widetilde{R}(\cdot)+p\mathrm{CNOT}(\cdot)\mathrm{CNOT}^{\dagger}$, respectively, for $p\in[0,1]$. The von Neumann conditional entropies of $\mathcal{N}^{p,\mathrm{SWAP}(2)}_{A^{\prime}B^{\prime}\to AB}$ and $\mathcal{N}^{p,\mathrm{CNOT}}_{A^{\prime}B^{\prime}\to AB}$ are, see Figure 1,
$S[A|B]_{\mathcal{N}^{p,\mathrm{SWAP}(2)}}$ $=-\frac{15p}{16}\log\frac{p}{16}-\left(1-\frac{15p}{16}\right)\log(1-\frac{15p}{16})-3,$ (75)
$S[A|B]_{\mathcal{N}^{p,\mathrm{CNOT}}}$ $=-\frac{15p}{16}\log\frac{p}{16}+\left(1-\frac{15p}{16}\right)\log(1-\frac{15p}{16})-\left(1-\frac{p}{2}\right)\log(\frac{1}{2}-\frac{p}{4})+\frac{p}{2}\log\frac{p}{4}-1.$ (76)